# On the Fundamental Limits of LLMs at Scale

### v1

@@ 3.4 Attention Computation Limits: Softmax Dynamics and Memory Complexity
Chunk boundaries break dependencies Errors at the window cuts; cross-chunk links lost. Bridges: memory tokens per chunk; summary vectors with attention into prior chunks; overlap windows with learned dedup; cached entity graphs. Dai et al. (2019); Wu et al. (2022); Sun and others (2024)
@@
External memory is not persistent across tasks/sessions Forgets earlier sessions; cannot keep project-scale facts. Persistent stores: vector DB + RAG with typed schemas; tool-augmented retrieval (program-of-thought); long-term key–value memory distilled from transcripts/logs. Lewis et al. (2020b); Schick et al. (2023)
@@
Evaluation misalignments (proxy tasks, short-context fine-tunes) Improvements don’t transfer to real long-doc reasoning. Benchmarks: use LongBench/NeedleBench/ $\infty$ Bench style tasks with dispersion & density controls; report accuracy vs. length, energy/token, and effective-context curves; ablate position reweighting. Bai and others (2023); Li and others (2024); Zhang and others (2024); Wang and others (2024)
@@
Alternatives to pure Transformers (state-space and hybrids) Attention saturation or cost dominates at scale. SSM class: S4/Hyena/Mamba for linear-time long-range processing; Hybrids: LongMamba (local attention + global state); gated recurrence for persistent memory; cross-architecture distillation from long-attention teachers. Gu and Dao (2023); Gu et al. (2021); Sun and others (2024)
@@ 4.1 Why Likelihood Falls Short
Current LLMs (e.g., GPT-3/ChatGPT/GPT-4, PaLM 2, InstructGPT, LaMDA) function like a ‘Mad Libs’ game: they fill in blanks using statistical associations rather than understanding and logically manipulating facts (Zhang, 2024). The probabilistic modeling approach captures correlations in language but does not guarantee that the model has learned the underlying functions or rules needed to solve reasoning problems Zhao et al. (2024). Many reasoning tasks (arithmetic, logic puzzles, algorithmic procedures, etc.) require precise intermediate computations and compositional generalization, the ability to combine learned components in novel ways. Models learn the pieces but often fail to compose them: in Boolean logic, accuracy drops sharply as expressions grow deeper or wider, even when the basic primitives are mastered. This suggests they rely on surface patterns rather than underlying rules (Kim and Thorne, ).
@@ 4.2 What Causes Reasoning to Fail in LLMs?
At the meta-level, models must learn when extended reasoning is warranted. Let $c\in\{\texttt{short},\texttt{think}\}$ be a control token; Thinkless, Fang et al. (2025) optimizes
@@
reducing unnecessary long-CoTs while preserving accuracy (Fang et al., 2025). Thus, effective systems guide search toward promising partial derivations, explicitly trade off accuracy against computational expense, and route adaptively between terse and deliberative modes, avoiding both premature commitment and unproductive overthinking (Zhang et al., 2024e; Sel et al., ). A related approach is hierarchical reasoning, where a model first picks abstract subgoals and then expands them into concrete steps (e.g., Hierarchical Reasoning Model (HRM) Wang et al. (2025a), HyperTree Planning (HTP) Gui et al. (2025)). But these methods often falter: errors in planning cascade downward, high-level plans may misalign with executable detail, and the space of possible subplans still grows rapidly unless heavily pruned. Thus, HRMs need subgoal verification, pruning, and cost awareness to be reliable (as discussed in Section 4.3).
@@ 4.3.1 Instantiate Reasoning using the Unified Objective
$\mathcal{R}_{\text{prompt}}(P,Q)=\text{LLM}\left(P\oplus Q\oplus\mathcal{P}_{\text{CoT}}\right)\text{ or }\underset{\text{path}}{\text{Search}}\left[\text{ToT}(P,Q)\right],$ (38)
@@
Fine-tuning modifies the model parameters to incorporate logical reasoning directly. This targets gaps in logic-rich supervision by adding rules, proofs, and structured derivations, particularly logical multi-step deduction and proofs, in pre-training corpora composed mainly of human-written texts that exhibit reflexive thinking rather than rigid logical reasoning (Morishita et al., 2024). The framework adaptation for fine-tuning methods (as shown in Figure 6) becomes:
@@ 4.3.2 Reasoning Consistency within the Unified Objective
which resembles the Grandor Chase decoder paradigm in information theory, where bits are iteratively flipped until all parity checks are satisfied. Here, logical constraints $\mathcal{C}_{ij}$ play an analogous role to parity relations, and consistency enforcement acts as a reasoning-parity check that iteratively adjusts answers until the collective chain of reasoning achieves global coherence. In this view, parity in AI reasoning corresponds to maintaining causal and logical alignment across multiple inferential steps.
@@
Figure 7: Figure: Overview of reasoning strategies, (a) Chain-of-Verification (generate, execute, verify), (b) Program-Aided Language (code-based execution), (c) Neuro-Symbolic Integration (neural + rules for multi-step logic), and (d) Multi-modal LLM (image+language grounding).
@@ 5 Why Does Retrieval Fail?
Retrieval-Augmented Generation (RAG) has emerged as a critical paradigm for enhancing Large Language Models (LLMs) by integrating external knowledge sources Lewis et al. (2020a). However, the performance of LLMs in RAG settings is heavily contingent upon the quality and relevance of the retrieved information Gupta et al. (2024). Formally, the retrieved passages condition the model’s solution space by acting as contextual evidence during generation:
$P(y\mid x,D_{r})=\int P(y\mid x,z)\,P(z\mid x,D_{r})\,dz$
@@
Two fundamental dimensions have been central to the research community’s efforts in advancing Retrieval-Augmented Generation (RAG) systems: (1) improving the quality of retrieval, and (2) enhancing LLM robustness to misleading or contradictory retrievals. These two axes capture the core limitations that dictate the overall reliability and factual consistency of RAG pipelines. In this section, we primarily focus on these dimensions. First, we discuss challenges related to retrieval quality i.e., how the precision, recall, and contextual alignment of retrieved documents influence generative performance. Second, robustness of LLMs to misretrieved or contradictory information, where retrieval errors can mislead the model’s reasoning trajectory or bias its token-level generation toward incorrect conclusions.
@@ 5.1.1 Relevance–Coverage Dilemma
Recent works address this balance through structure-aware retrieval, where compact graph or cluster representations encode semantically related content within fewer tokens Cheng et al. (2025). Graph-based retrievers leverage node-level relevance propagation to preserve contextual diversity under budget constraints Zhu et al. (2025a); Guo et al. (2025b). Similarly, KG-guided and hierarchical retrieval methods compress correlated information before injection Li et al. (2024d). While such designs partially alleviate redundancy, inherent bottlenecks persist: incomplete graph connectivity limits factual recall, graph linearization reintroduces token overhead, and traversal-based ranking adds computational latency. Hence, the relevance–coverage dilemma remains a structural limitation of token-bounded retrieval pipelines.
@@ 5.1.2 Information Discretization in Token-Bounded Retrieval
This optimization assumes chunk independence, yet most natural language dependencies are cross-chunk. When semantically linked units (e.g., condition–action pairs) straddle boundaries, $s(q,c_{i})$ and $s(q,c_{i+1})$ both drop below threshold, eliminating essential context. The resulting fragmentation loss causes retrieval to return incomplete evidence, which the generator then over-interprets as complete input, degrading factual precision and calibration Qian et al. (2024); Lu et al. (2025).
Hierarchical and chunk-free retrievals alleviate but do not eliminate this degradation. Hierarchical chunking dynamically merges semantically adjacent spans to preserve local coherence Lu et al. (2025), while chunk-free in-context retrieval embeds full documents and extracts spans directly Qian et al. (2024); Brådland et al. (2025). Both improve retrieval fidelity yet remain bounded by the token limit $B$: even if the retriever identifies coherent evidence, it must still serialize it into a finite context, discarding residual dependencies.
Beyond fragmentation, relevance degradation arises from ranker-level phenomena. As the retriever expands $D_{r}$ to increase coverage, average similarity $\bar{s}(q,D_{r})=\frac{1}{|D_{r}|}\sum_{c\in D_{r}}s(q,c)$ decreases, reducing signal-to-noise in the prompt. Document-level mismatches and embedding sensitivity further compound error, as minor paraphrases can alter embedding neighborhoods, leading to unstable recall@k and inconsistent grounding Cao et al. (2025); Park and Lee (2024).
@@
yielding an exposure likelihood reflected by the retrieval metric $\mathrm{Recall@}k=\frac{|D_{r}\cap S^{\star}|}{|S^{\star}|}$. Whenever $\exists\,d^{\star}\in S^{\star}$ with rank >k>k, the model’s generation is upper-bounded by missing evidence, regardless of reasoning capacity. This manifests as the scattered-evidence failure, where the rationale spans multiple documents but only a subset appears in $D_{r}$, fragmenting multi-hop inference. Even multi-passage fusion techniques (e.g., FiD-style decoding) remain constrained by ranker exposure and token budgets Izacard and Grave (2020); Agrawal et al. (2024).
Positional or order bias further compounds this issue. Long-context analyses reveal a consistent “lost-in-the-middle” effect: tokens located near the start and end of the prompt receive disproportionately higher attention weights than mid-sequence tokens. Empirically, relocating gold passages from the extremes to mid-context reduces answer recall, confirming a primacy–recency weighting curve $w(t)$ with $w(0),w(1)\gg w(0.5)$ Liu et al. (2023c); Li et al. (2024b). This bias arises from causal masking and positional encoding decay in Transformer layers—RoPE-based encodings exhibit long-range attenuation that weakens middle-context dependency modeling Kazemnejad et al. (2023). Consequently, document ordering within the retrieved set materially affects output; reordering equally scored passages alters factual accuracy, demonstrating the model’s sensitivity to serial position Cuconasu et al. (2025); Hsieh et al. (2024).
@@
illustrating that generation fidelity depends jointly on exposure (ranking) and utilization (attention). While reranking, iterative decoding, and attention-based reordering alleviate some degradation Reddy et al. (2024); Li et al. (2024b), the bottleneck persists: the serialized, token-bounded prompt enforces both top-$k$ truncation and position-dependent conditioning. Once critical evidence is omitted or placed unfavorably within the context, the posterior $P(y\!\mid\!x,D_{r})$ cannot recover it. Retrieval quality in RAG is therefore limited not only by what is retrieved, but by where and how it is positioned within the LLM’s representational scope.
@@
RAG systems inherit a critical vulnerability from their external memory: any retriever that ranks based on similarity $s(q,d)=\langle f_{Q}(q),f_{D}(d)\rangle$ can be adversarially biased through minimal perturbations of the knowledge base Xian et al. (2024); Zou et al. (2025). Let the retrieval corpus be $\mathcal{D}=\mathcal{D}_{\text{clean}}\cup\mathcal{P}$, where $\mathcal{P}$ is a small poisoned subset. The retriever returns
@@
subject to a small injection budget $|\mathcal{P}|=N\ll|\mathcal{D}_{\text{clean}}|$. Empirically, PoisonedRAG demonstrates that inserting only five poisoned documents per target query in million-scale KBs achieves $\sim$90% attack success rates, establishing a highly efficient adversarial regime Zou et al. (2025).
@@
i.e., they remain orthogonal to clean evidence yet maximally aligned with the query embedding. This displaces authentic evidence from the top-$k$ frontier under the same token budget $B$, corrupting retrieval without overtly altering corpus semantics. Because retrievers reuse the same embedding manifold for paraphrased queries $q^{\prime}\!\sim\!\mathcal{Q}(q^{\star})$, poisons generalize across lexical variants, producing a knowledge corruption cascade Chang et al. (2025); Ha et al. (2025):
@@
Activation-level detectors such as RevPRAG estimate a poisoning indicator $z=\sigma(g(h_{1:L}))$ from intermediate LLM activations $h_{1:L}$, achieving $\sim$98% true-positive rate with $2\%$ false positives across retrieval architectures Tan et al. (2024); Zhang et al. (2025a). Forensic traceback methods model removal-based counterfactuals—finding the minimal set $\widehat{\mathcal{P}}\subset\mathcal{D}$ whose exclusion flips the generation outcome, i.e.,

### v2

@@ 3.4 Attention Computation Limits: Softmax Dynamics and Memory Complexity
Chunk boundaries break dependencies Errors at the window cuts; cross-chunk links lost. Bridges: memory tokens per chunk; summary vectors with attention into prior chunks; overlap windows with learned dedup; cached entity graphs. (Dai et al., 2019; Wu et al., 2022; Sun and others, 2024)
@@
External memory is not persistent across tasks/sessions Forgets earlier sessions; cannot keep project-scale facts. Persistent stores: vector DB + RAG with typed schemas; tool-augmented retrieval (program-of-thought); long-term key–value memory distilled from transcripts/logs. (Lewis et al., 2020b; Schick et al., 2023)
@@
Evaluation misalignments (proxy tasks, short-context fine-tunes) Improvements don’t transfer to real long-doc reasoning. Benchmarks: use LongBench/NeedleBench/ $\infty$ Bench style tasks with dispersion & density controls; report accuracy vs. length, energy/token, and effective-context curves; ablate position reweighting. (Bai et al., 2024b; Li and others, 2024; Zhang and others, 2024; Wang and others, 2024)
@@
Alternatives to pure Transformers (state-space and hybrids) Attention saturation or cost dominates at scale. SSM class: S4/Hyena/Mamba for linear-time long-range processing; Hybrids: LongMamba (local attention + global state); gated recurrence for persistent memory; cross-architecture distillation from long-attention teachers. (Gu and Dao, 2023; Gu et al., 2021; Sun and others, 2024)
@@ 4.1 Why Likelihood Falls Short
Current LLMs (e.g., GPT-3/ChatGPT/GPT-4, PaLM 2, InstructGPT, LaMDA) function like a ‘Mad Libs’ game: they fill in blanks using statistical associations rather than understanding and logically manipulating facts (Zhang, 2024). The probabilistic modeling approach captures correlations in language but does not guarantee that the model has learned the underlying functions or rules needed to solve reasoning problems (Zhao et al., 2024). Many reasoning tasks (arithmetic, logic puzzles, algorithmic procedures, etc.) require precise intermediate computations and compositional generalization, the ability to combine learned components in novel ways. Models learn the pieces but often fail to compose them: in Boolean logic, accuracy drops sharply as expressions grow deeper or wider, even when the basic primitives are mastered. This suggests they rely on surface patterns rather than underlying rules (Kim and Thorne, 2024).
@@ 4.2 What Causes Reasoning to Fail in LLMs?
At the meta-level, models must learn when extended reasoning is warranted. Let $c\in\{\texttt{short},\texttt{think}\}$ be a control token; Thinkless, (Fang et al., 2025) optimizes
@@
reducing unnecessary long-CoTs while preserving accuracy (Fang et al., 2025). Thus, effective systems guide search toward promising partial derivations, explicitly trade off accuracy against computational expense, and route adaptively between terse and deliberative modes, avoiding both premature commitment and unproductive overthinking (Zhang et al., 2024e; Sel et al., ). A related approach is hierarchical reasoning, where a model first picks abstract subgoals and then expands them into concrete steps (e.g., Hierarchical Reasoning Model (HRM) (Wang et al., 2025a), HyperTree Planning (HTP) (Gui et al., 2025)). But these methods often falter: errors in planning cascade downward, high-level plans may misalign with executable detail, and the space of possible subplans still grows rapidly unless heavily pruned. Thus, HRMs need subgoal verification, pruning, and cost awareness to be reliable (as discussed in Section 4.3).
@@ 4.3.1 Instantiate Reasoning using the Unified Objective
$\mathcal{R}_{\text{prompt}}(P,Q)=\text{LLM}\left(P\oplus Q\oplus\mathcal{P}_{\text{CoT}}\right)\text{ or }\underset{\text{path}}{\text{Search}}\left[\text{ToT}(P,Q)\right].$ (38)
@@
Fine-tuning modifies model parameters to incorporate logical reasoning directly. This targets gaps in logic-rich supervision by adding rules, proofs, and structured derivations, particularly logical multi-step deduction and proofs, in pre-training corpora composed mainly of human-written texts that exhibit reflexive thinking rather than rigid logical reasoning (Morishita et al., 2024). The framework adaptation for fine-tuning methods (as shown in Figure 6) becomes:
@@ 4.3.2 Reasoning Consistency within the Unified Objective
This resembles the Grandor Chase decoder paradigm in information theory, where bits are iteratively flipped until all parity checks are satisfied. Here, logical constraints $\mathcal{C}_{ij}$ play an analogous role to parity relations, and consistency enforcement acts as a reasoning-parity check that iteratively adjusts answers until the collective chain of reasoning achieves global coherence. In this view, parity in AI reasoning corresponds to maintaining causal and logical alignment across multiple inferential steps.
@@
Figure 7: Overview of reasoning strategies, (a) Chain-of-Verification (generate, execute, verify), (b) Program-Aided Language (code-based execution), (c) Neuro-Symbolic Integration (neural + rules for multi-step logic), and (d) Multi-modal LLM (image+language grounding).
@@ 5 Why Does Retrieval Fail?
Retrieval-Augmented Generation (RAG) has emerged as a critical paradigm for enhancing Large Language Models (LLMs) by integrating external knowledge sources (Lewis et al., 2020a). However, the performance of LLMs in RAG settings is heavily contingent upon the quality and relevance of the retrieved information (Gupta et al., 2024). Formally, the retrieved passages condition the model’s solution space by acting as contextual evidence during generation:
$P(y\mid x,D_{r})=\int P(y\mid x,z)\,P(z\mid x,D_{r})\,dz.$
@@
Two fundamental dimensions have been central to the research community’s efforts in advancing Retrieval-Augmented Generation (RAG) systems: (1) improving the quality of retrieval, and (2) enhancing LLM robustness to misleading or contradictory retrievals. These two axes capture the core limitations that dictate the overall reliability and factual consistency of RAG pipelines. In this section, we primarily focus on these dimensions. First, we discuss challenges related to retrieval quality, i.e., how the precision, recall, and contextual alignment of retrieved documents influence generative performance. Second, robustness of LLMs to misretrieved or contradictory information, where retrieval errors can mislead the model’s reasoning trajectory or bias its token-level generation toward incorrect conclusions.
@@ 5.1.1 Relevance–Coverage Dilemma
Recent works address this balance through structure-aware retrieval, where compact graph or cluster representations encode semantically related content within fewer tokens (Cheng et al., 2025). Graph-based retrievers leverage node-level relevance propagation to preserve contextual diversity under budget constraints (Zhu et al., 2025a; Guo et al., 2025b). Similarly, KG-guided and hierarchical retrieval methods compress correlated information before injection (Li et al., 2024d). While such designs partially alleviate redundancy, inherent bottlenecks persist: incomplete graph connectivity limits factual recall, graph linearization reintroduces token overhead, and traversal-based ranking adds computational latency. Hence, the relevance–coverage dilemma remains a structural limitation of token-bounded retrieval pipelines.
@@ 5.1.2 Information Discretization in Token-Bounded Retrieval
This optimization assumes chunk independence, yet most natural language dependencies are cross-chunk. When semantically linked units (e.g., condition–action pairs) straddle boundaries, $s(q,c_{i})$ and $s(q,c_{i+1})$ both drop below threshold, eliminating essential context. The resulting fragmentation loss causes retrieval to return incomplete evidence, which the generator then over-interprets as complete input, degrading factual precision and calibration (Qian et al., 2024; Lu et al., 2025).
Hierarchical and chunk-free retrievals alleviate but do not eliminate this degradation. Hierarchical chunking dynamically merges semantically adjacent spans to preserve local coherence (Lu et al., 2025), while chunk-free in-context retrieval embeds full documents and extracts spans directly (Qian et al., 2024; Brådland et al., 2025). Both improve retrieval fidelity yet remain bounded by the token limit $B$: even if the retriever identifies coherent evidence, it must still serialize it into a finite context, discarding residual dependencies.
Beyond fragmentation, relevance degradation arises from ranker-level phenomena. As the retriever expands $D_{r}$ to increase coverage, average similarity $\bar{s}(q,D_{r})=\frac{1}{|D_{r}|}\sum_{c\in D_{r}}s(q,c)$ decreases, reducing signal-to-noise in the prompt. Document-level mismatches and embedding sensitivity further compound error, as minor paraphrases can alter embedding neighborhoods, leading to unstable recall@k and inconsistent grounding (Cao et al., 2025; Park and Lee, 2024).
@@
yielding an exposure likelihood reflected by the retrieval metric $\mathrm{Recall@}k=\frac{|D_{r}\cap S^{\star}|}{|S^{\star}|}$. Whenever $\exists\,d^{\star}\in S^{\star}$ with rank >k>k, the model’s generation is upper-bounded by missing evidence, regardless of reasoning capacity. This manifests as the scattered-evidence failure, where the rationale spans multiple documents but only a subset appears in $D_{r}$, fragmenting multi-hop inference. Even multi-passage fusion techniques (e.g., FiD-style decoding) remain constrained by ranker exposure and token budgets (Izacard and Grave, 2020; Agrawal et al., 2024).
Positional or order bias further compounds this issue. Long-context analyses reveal a consistent “lost-in-the-middle” effect: tokens located near the start and end of the prompt receive disproportionately higher attention weights than mid-sequence tokens. Empirically, relocating gold passages from the extremes to mid-context reduces answer recall, confirming a primacy–recency weighting curve $w(t)$ with $w(0),w(1)\gg w(0.5)$ (Liu et al., 2023c; Li et al., 2024b). This bias arises from causal masking and positional encoding decay in Transformer layers—RoPE-based encodings exhibit long-range attenuation that weakens middle-context dependency modeling (Kazemnejad et al., 2023). Consequently, document ordering within the retrieved set materially affects output; reordering equally scored passages alters factual accuracy, demonstrating the model’s sensitivity to serial position (Cuconasu et al., 2025; Hsieh et al., 2024).
@@
illustrating that generation fidelity depends jointly on exposure (ranking) and utilization (attention). While reranking, iterative decoding, and attention-based reordering alleviate some degradation (Reddy et al., 2024; Li et al., 2024b), the bottleneck persists: the serialized, token-bounded prompt enforces both top-$k$ truncation and position-dependent conditioning. Once critical evidence is omitted or placed unfavorably within the context, the posterior $P(y\!\mid\!x,D_{r})$ cannot recover it. Retrieval quality in RAG is therefore limited not only by what is retrieved, but by where and how it is positioned within the LLM’s representational scope.
@@
RAG systems inherit a critical vulnerability from their external memory: any retriever that ranks based on similarity $s(q,d)=\langle f_{Q}(q),f_{D}(d)\rangle$ can be adversarially biased through minimal perturbations of the knowledge base (Xian et al., 2024; Zou et al., 2025). Let the retrieval corpus be $\mathcal{D}=\mathcal{D}_{\text{clean}}\cup\mathcal{P}$, where $\mathcal{P}$ is a small poisoned subset. The retriever returns
@@
subject to a small injection budget $|\mathcal{P}|=N\ll|\mathcal{D}_{\text{clean}}|$. Empirically, PoisonedRAG demonstrates that inserting only five poisoned documents per target query in million-scale KBs achieves $\sim$90% attack success rates, establishing a highly efficient adversarial regime (Zou et al., 2025).
@@
i.e., they remain orthogonal to clean evidence yet maximally aligned with the query embedding. This displaces authentic evidence from the top-$k$ frontier under the same token budget $B$, corrupting retrieval without overtly altering corpus semantics. Because retrievers reuse the same embedding manifold for paraphrased queries $q^{\prime}\!\sim\!\mathcal{Q}(q^{\star})$, poisons generalize across lexical variants, producing a knowledge corruption cascade (Chang et al., 2025; Ha et al., 2025):
@@
Activation-level detectors such as RevPRAG estimate a poisoning indicator $z=\sigma(g(h_{1:L}))$ from intermediate LLM activations $h_{1:L}$, achieving $\sim$98% true-positive rate with $2\%$ false positives across retrieval architectures (Tan et al., 2024; Zhang et al., 2025a). Forensic traceback methods model removal-based counterfactuals—finding the minimal set $\widehat{\mathcal{P}}\subset\mathcal{D}$ whose exclusion flips the generation outcome, i.e.,